# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v2

@@ Appendix D Hyperpipeline examples
In Section II.7 we presented our iterative strategy for hyperparameter inference. This approach has been used for GW source parameter inference with RIFT as well as the equation of state which are presented in the Section III. Here, we provide illustrations showing how our iterative strategy works within the context of EoS hyperparameter inference.
@@ Appendix E NICER PSR-J0437
One of the NICER collaboration groups announced the observational constraints for PSR-J0437 [142]. The pulsar has a strong a-priori mass constraint from radio observations [180]. The full posterior data release has not occurred, however, here we apply this pulsar’s separate 90% credible intervals for its mass and radius and can construct an approximate Gaussian posterior to explore its effect on the EoS. The reported Radius is $R=11.36^{+0.95}_{-0.63}~{}{\rm km}$ and $M=1.418\pm 0.037~{}{\rm M_{\odot}}$. For this preliminary application, we employ the constraints as a Gaussian with $R=11.36\pm 0.95~{}{\rm km}$ and $M=1.418\pm 0.037~{}{\rm M_{\odot}}$.

### v3

@@ Appendix D Hyperpipeline examples
In Section 2.9 we presented our iterative strategy for hyperparameter inference. This approach has been used for GW source parameter inference with RIFT as well as the equation of state which are presented in the Section 3. Here, we provide illustrations showing how our iterative strategy works within the context of EoS hyperparameter inference.
@@ Appendix E NICER PSR-J0437
One of the NICER collaboration groups announced the observational constraints for PSR-J0437 (2024ApJ...971L..20C). The pulsar has a strong a-priori mass constraint from radio observations (2023PASA...40...49Z). The full posterior data release has not occurred, however, here we apply this pulsar’s separate 90% credible intervals for its mass and radius and can construct an approximate Gaussian posterior to explore its effect on the EoS. The reported Radius is $R=11.36^{+0.95}_{-0.63}\penalty 10000\ {\rm km}$ and $M=1.418\pm 0.037\penalty 10000\ {\rm M_{\odot}}$. For this preliminary application, we employ the constraints as a Gaussian with $R=11.36\pm 0.95\penalty 10000\ {\rm km}$ and $M=1.418\pm 0.037\penalty 10000\ {\rm M_{\odot}}$.